# Erased, But Not Forgotten: Erased Rectified Flow Transformers Still Remain Unsafe Under Concept Attack

### v2

@@ 4 Method
Overview. Section 3 shows that amplifying target tokens drives the attention maps into divergence, yielding collapsed or distorted images. Our goal is therefore precise: restore the erased concept while preventing attention divergence and preserving global layout and unrelated content. To this end, we introduce ReFlux, a fine-tuning strategy that employs LoRA on Flux’s text parameters.
Attention reactivation loss. The first step is to stabilize convergence. Direct amplification of erased tokens tends to destabilize optimization, so our objective is to reactivate suppressed attention while ensuring stable training dynamics. In our setting, the optimization variable is the attention weight $\mathbf{W}_{\mathtt{attn}}[:,:,i]$ for each target index $i\in\mathtt{target\_idx}$, and the driving term is $\sum_{i\in\mathtt{target\_idx}}\mathbf{W}_{\mathtt{attn}}[:,:,i]$. Following the principle of proximal optimization (Parikh & Boyd, 2013), we formulate the update as a trust-region step that maximizes the gain on target attention while penalizing deviations from the pretrained distribution, ensuring both progress and stability:
$\mathbf{W}^{+}_{\mathtt{attn}}=\arg\max_{\mathbf{W}_{\mathtt{attn}}}\Big\{\,\sum_{i\in\text{target\_idx}}\mathbf{W}_{\mathtt{attn}}[:,:,i]\;-\;\frac{1}{2\lambda_{d}}\,D(\mathbf{W}_{\mathtt{attn}},\mathbf{W}^{(0)}_{\mathtt{attn}})\,\Big\}$, (3)
where $\mathbf{W}^{(0)}_{\mathtt{attn}}$ denotes the baseline attention map and $\lambda_{d}>0$ controls the step size. $D(\cdot,\cdot)$ denotes a compound distance penalty that jointly constrains step size and preserves distributional stability:
$D(\mathbf{W}_{\mathtt{attn}},\mathbf{W}^{(0)}_{\mathtt{attn}})=\lambda\|\mathbf{W}_{\mathtt{attn}}-\mathbf{W}^{(0)}_{\mathtt{attn}}\|_{2}^{2}+2\tau\,\mathrm{KL}\!\Big(\mathtt{softmax}(\mathbf{W}_{\mathtt{attn}})\,\big\|\,\mathtt{softmax}(\mathbf{W}^{(0)}_{\mathtt{attn}})\Big)$, (4)
where $\|\mathbf{W}_{\mathtt{attn}}-\mathbf{W}^{(0)}_{\mathtt{attn}}\|_{2}^{2}$ penalizes updates that deviate excessively from the baseline, ensuring the optimization does not stray away from the target concept, while the KL term enforces the updated attention distribution to remain aligned with the original direction, thereby preserving the global layout and visual style (as visualized in Figure 6).
Expanding the KL divergence, with $p=\mathtt{softmax}(\mathbf{W}_{\mathtt{attn}})$ and $p_{0}=\mathtt{softmax}(\mathbf{W}^{(0)}_{\mathtt{attn}})$:
@@
where $H(p)=-\sum_{i}p_{i}\log p_{i}$ is the entropy and by choosing $p_{0}$ as the uniform distribution (an unbiased baseline), the last term becomes constant and can be omitted in loss. Substituting Equation 5 into the compound distance penalty in Equation 4, we obtain the final attention reactivation loss:
@@
This directly addresses the core objective of stabilizing convergence: it reactivates suppressed attention on target tokens while constraining updates to remain within a stable regime. Further implementation details are provided in Appendix E.
Attack-guided velocity loss. While the attention reactivation loss stabilizes convergence and restores signal strength at token level, it alone does not guarantee that the recovered attention will consistently manifest in generated images. To enhance robustness, we then complement it with a velocity-based objective that directly shapes the flow matching generative dynamics of Flux:
@@
$\mathcal{L}_{\text{attack}}=\mathbb{E}\Big[-\,v_{\theta+\Delta\theta}(x_{t},c_{\text{target}},t)\;-\;\eta\,\big\|v_{\theta}(x_{t},c_{\text{target}},t)-v_{\theta}(x_{t},\varnothing,t)\big\|_{2}^{2}\Big],$ (7)
where $x_{t}$ denotes the latent at timestep $t$ in the denoising process, $\theta$ are the original parameters, $\Delta\theta$ is the learnable LoRA weights for fine-tuning, and $\eta$ controls the aggressiveness of the attack, equivalent with the negative guidance in Gandikota et al. (2023). The conditional input $c_{\text{target}}$ specifies the concept intended for attack (e.g., “nudity”), while $\varnothing$ denotes the unconditional null prompt. Additionally, The term $v$ represents the velocity of the flow matching process, which forms the core of Flux’s scheduling mechanism and is conceptually equivalent to the v-prediction (Salimans & Ho, 2022) in diffusion models. By targeting this variable, our attack alters the generative trajectory, steering it away from suppression and restoring erased concepts with higher precision.
@@
LoRA consistency loss. While the previous objectives focus on reactivating suppressed attention and amplifying target velocity, they may inadvertently affect unrelated visual elements. For example, given the prompt “a cartoon naked girl in a hotel”, where the erased concept is “naked” and the unrelated content includes “cartoon” and “hotel”, our goal is to restore the concept “naked” without degrading the generation of other elements or overall visual layout. To achieve this, we fix a prompt $c$ that contains both the target and unrelated content, and sample 4–8 reference images using random seeds, then train the LoRA weights $\Delta\theta$ to minimally deviate from the original generation dynamics:
$\mathcal{L}_{\text{lora}}=\mathbb{E}_{(u_{\text{pix}},x_{T},t,c)\sim\mathcal{I}_{f}}\Big\|\,v-v_{\theta+\Delta\theta}(u_{t},c,t)\,\Big\|_{2}^{2},$ (8)
@@
where $v=x_{T}-u_{\text{pix}}$ is the ground-truth velocity ($T$ is the total denoising timesteps, $x_{T}$ is random noise, $u_{\text{pix}}$ is the VAE (Kingma, 2013) encoded latent of the sampled image from $\mathcal{I}_{f}$), and the noised latent is given by $u_{t}=(1-t)\,u_{\text{pix}}+t\,x_{T}$. This formulation directly follows the flow matching schedule in Flux, ensuring our attack is achieved without disturbing the global synthesis trajectory or unrelated content.
@@
Together, these three objectives form a unified and principled framework for concept attack. The attention reactivation loss revives suppressed attention signals, the attack-guided velocity loss amplifies flow matching dynamics to reinforce the concept, and the LoRA consistency loss preserves style fidelity while shielding unrelated content. By balancing precision, stability, and stealth, our method achieves robust and targeted reactivation of erased concepts in Flux.
@@ 5.1 Implementation Details
Setup. We conduct all experiments on the Flux.1 [dev] [^1] https://huggingface.co/black-forest-labs/FLUX.1-dev model, an openly released distilled rectified flow transformer that retains strong prompt fidelity and high generation quality, employing the flow-matching Euler sampler with 28 denoising steps. Our attack method is optimized for 1,000 steps using the AdamW optimizer (Loshchilov et al., 2017) with a learning rate of 0.001 and the negative guidance factor $\eta=1$. Following Gao et al. (2025), we update 3.57MB text-related parameters ($\mathtt{add\_q\_proj}$ and $\mathtt{add\_k\_proj}$) within the dual-stream blocks. All experiments are conducted on a single NVIDIA H20 GPU (96 GB VRAM) with batch size 1.
Baselines. We attack a suite of concept erasure strategies and compare our method against state-of-the-art attack baselines. For concept attack, we include the white-box methods UnlearnDiffAtk (Zhang et al., 2024) and P4D (Chin et al., 2024), the black-box methods Ring-A-Bell and Ring-A-Bell Union (Tsai et al., 2024), and the LLM reasoning-driven Reason2Attack (Zhang et al., 2025). For concept erasure, we evaluate against leading concept removal methods including ESD (Gandikota et al., 2023), AC (Kumari et al., 2023), EAP (Bui et al., 2024), EA (Gao et al., 2025) and CP (Chavhan et al., 2025). More implementation details are provided in Appendix F.
@@ 5.2 Results
Evaluation on NSFW concepts. NSFW concepts are widely recognized benchmarks. We evaluate our method using the I2P dataset (Schramowski et al., 2023), focusing on nudity and violence. For nudity, we select 109 prompts with more than 50% nudity percentage, with detection by NudeNet (Bedapudi, 2019) at a threshold of 0.6. For violence, following Tsai et al. (2024), we choose 235 prompts with more than 50% inappropriateness but less than 50% nudity and labeled as harmful, and apply the Q16-classifier (Schramowski et al., 2022) to detect harmful subjects. Finally, results are reported by attack success rate (ASR) following Zhang et al. (2024).
Table 2: Attack evaluation on NSFW concept benchmarks. We compare concept erasure methods against multiple attack baselines. Results are reported in terms of Attack Success Rate (ASR, %; lower is better), measured on the I2P benchmark. “No Attack” indicates absence of any attack method. The performance of the original Flux.1 [dev] model is shown as a reference.
Concept Methods Flux.1 AC ESD-1 ESD-3 EA EAP CP
No Attack 44.04 32.11 21.10 14.67 29.35 44.04 10.09
Ring-A-Bell 40.36 27.50 11.01 18.34 29.35 35.77 11.92
Ring-A-Bell Union 65.14 44.03 24.77 29.36 42.20 45.87 20.18
Reason2Attack 67.88 73.39 59.63 56.88 71.55 69.72 40.36
UnlearnDiffAtk 100.00 85.32 76.14 70.64 82.57 98.16 64.22
@@
No Attack 65.53 64.68 54.89 58.72 53.19 61.28 48.51
Ring-A-Bell 75.74 74.89 66.80 68.93 51.48 65.53 50.21
Ring-A-Bell Union 81.70 80.00 77.02 82.55 61.28 79.57 52.76
Reason2Attack 71.91 67.23 61.28 61.70 57.44 70.63 41.70
UnlearnDiffAtk 87.23 85.10 80.85 84.25 78.72 86.38 55.74
@@
Table 2 presents our results for attacking prevailing erasing methods. Our attack achieves the highest ASR across all settings, consistently outperforming existing baselines. Notably, Ring-A-Bell fails to mount effective attacks on Flux as analyzed in Section 3. Among erasure methods, CP (Chavhan et al., 2025) proves to be a robust strategy on Flux. In contrast, methods that traditionally optimize cross-attention layers (e.g., AC and ESD) exhibit reduced effectiveness when transferred from the U-Net design of SD to the transformer architecture of Flux. This limitation is known as concept residue, i.e., the incomplete removal of concepts, making them prone to reactivation and even amplification. As shown in Figure 3, our method restores erased concepts with high fidelity, such as exposed body details in nudity cases and blood patterns in violence, while preserving consistency of unrelated elements (e.g., hairstyle, posture, background layout), whereas other attacks often fail to sustain success or produce substantial visual changes.
Evaluation on artistic styles. Here, we evaluate on two recognized painting styles: Van Gogh and Pablo Picasso, using the dataset of Chavhan et al. (2025) that provides 50 prompts per style. Following Zhang et al. (2024), we finetune an ImageNet-pretrained ViT-base model on WikiArt (Saleh & Elgammal, 2015) and obtain a 129-class style classifier for evaluation. In reporting ASR, to avoid overly restrictive, both Top-1 and Top-3 ASR are presented, depending on whether the result is classified as the target style as the top prediction or within the top three.
Table 3 shows our attack outperforms all baselines in 9 out of 12 evaluation settings, demonstrating clear and consistent improvements across artistic style benchmarks. This systematic advantage shows the effectiveness of attention reactivation and velocity control for global and highly abstract concepts, while attacks that optimize prompt embeddings are unstable and less effective on Flux.
@@
Figure 3: Visual comparison of nudity and violence across attack methods under different erasure strategies. Yellow framed images are original generations from Flux.1 [dev]. Blue bars and blurring are added for publication purposes. Our method achieves precise concept reactivation while preserving the layout and unrelated elements of the images.
Figure 4: Visual results on Entity, Abstraction, and Relationship. The leftmost column shows the defense effect of EA. Yellow framed images denote the original generations from Flux.1 [dev].
Evaluation on miscellaneousness. We evaluate our method on 3 broader categories: Entity, Abstraction and Relationship. Here, we choose 10 concepts for each category (see Appendix F for full list of concepts) and adopt CLIP classification as the measuring metrics. As shown in Table 4, our attack consistently outperforms baselines, with stronger gains on abstract and relational concepts, confirming its stable effectiveness across both concrete and abstract categories.
Table 4: Attack evaluation on specific category benchmarks: Entity (e.g., soccer, car), Abstraction (e.g., green, two) and Relationship (e.g., hug, back to back). CLIP classification accuracies are reported for each category. All presented values are denoted in percentage (%).
@@
Figure 5: Layer-wise attack gains across Flux dual stream blocks. Attack gain is defined as the relative increase in activation strength of our attack method compared to the erased model (EA).
Figure 4 presents the visual comparison of our method and UnlearnDiffAtk when attacking the EA erasure strategy. It is evident that our approach remains robust even for broader and more abstract concepts, where UnlearnDiffAtk may fail or produce images of poor quality. Additionally, a key strength of our method lies in preserving the overall layout and retaining unrelated elements.
Evaluation on efficiency. Beyond the strong performance analyzed above, our attack also achieves notable efficiency. Attacking a single prompt requires only the time to generate one image (about 0.5 minutes), whereas UnlearnDiffAtk and P4D depend on a costly PGD process exceeding 20 minutes. This efficiency advantage reflects the lightweight design and practical applicability of our approach.
Evaluation on layout consistency. Table 5 reports average similarity between images before and after attacking EA, computed using a MobileNet (Howard et al., 2017)-based utility. Results show that our method balances high ASR with strong preservation of layout and unrelated content.
Evaluation on layer-wise gains. To probe how our attack engages Flux’s internal representations, we analyze the 19 dual-stream blocks. As shown in Figure 5, erased concepts are restored at distinct stages: for “nudity”, strong gains emerge in early layers and reappear at the final block; for “Van Gogh”, clear peaks occur in both middle and late layers; and for “soccer”, improvements are more diffuse yet consistently positive. These patterns indicate that our attack exploits specific representation stages rather than acting uniformly across layers, with further analysis provided in Appendix G.
Table 5: Average image consistency (%) between erased (EA) and attacked outputs.
Method Nudity Violence Style Entity Abstract Relation
UnlearnDiffAtk 55.77 41.42 67.66 47.61 55.11 63.97
Ring-A-Bell Union 42.64 64.14 74.12 71.17 61.32 73.77
P4D - - 65.33 - - -
Reason2Attack 74.34 36.64 - - - -
@@
Ablation study. To evaluate the effectiveness of our loss functions, we conduct an ablation study on celebrities. We attack the EA erasure method using Gao et al. (2025) dataset (a subset of CelebA (Liu et al., 2018)), containing 50 celebrities for attack and 50 for retention.
Different variations are quantified in Table 6. The velocity objective $\mathcal{L}_{\text{attack}}$ provides the foundation of robustness by steering generative dynamics away from suppression, enhancing the reliability of concept reactivation across settings. The attention reactivation term $\mathcal{L}_{\text{attn}}$ sharpens targeting by directly amplifying suppressed attention signals, significantly improving Acce. As shown in Figure 6, its L2 regularizer constrains activation strength for reliable concept revival, and entropy maintains image style and visual coherence. Additionally, the consistency term $\mathcal{L}_{\text{lora}}$ further anchors background and irrelevant attributes, keeping details such as clothing, hairstyle, and posture consistent between the erased and attacked output, serving as a key contributor to high Accir. When integrated, these components form a synergistic objective: attention reactivation drives precision, velocity control enhances robustness, and LoRA regularization safeguards fidelity, together delivering the strongest and most balanced reactivation of erased concepts.
@@
Others. Additional details and results are provided in Appendix H, including further ablation with different fine-tuning parameters, the complete dataset used in our study, details of recognition classifiers, extended visualizations, user study and visual-language model assessments.

### v3

@@ 4 Method
As discussed in Section 3, simply boosting the attention on target tokens causes the maps to blow up, leading to distorted images. Our goal is straightforward: bring back the erased concept without letting attention diverge, while keeping the overall layout and unrelated content intact. To achieve this, we introduce ReFlux, a lightweight fine-tuning strategy that applies LoRA to Flux’s text parameters.
Attention reactivation loss. We start by stabilizing convergence. Directly boosting the erased tokens often makes optimization unstable, so our goal is to bring suppressed attention back while keeping the training dynamics steady. In our setup, the optimization variable is the attention weight $\mathbf{W}_{\mathtt{attn}}[:,:,i]$ for each target index $i\in\mathtt{target\_idx}$, and the driving term is $\sum{i\in\mathtt{target\_idx}}\mathbf{W}_{\mathtt{attn}}[:,:,i]$. Following the idea of proximal optimization [33], we frame the update as a trust-region step that increases target attention while penalizing large deviations from the pretrained distribution, achieving both steady progress and stability:
[…]
where $\mathbf{W}^{(0)}_{\mathtt{attn}}$ is the baseline attention map and $\lambda_{d}>0$ controls the step size. $D(\cdot,\cdot)$ represents a compound distance term that limits the update magnitude while keeping the overall attention distribution stable:
[…]
where $\|\mathbf{W}_{\mathtt{attn}}-\mathbf{W}^{(0)}_{\mathtt{attn}}\|_{2}^{2}$ penalizes large deviations from the baseline, keeping the optimization close to the target concept. The KL term encourages the updated attention distribution to stay aligned with the original direction, helping preserve overall layout and visual style (see Figure 6).
Expanding the KL divergence, where $p=\mathtt{softmax}(\mathbf{W}_{\mathtt{attn}})$ and $p_{0}=\mathtt{softmax}(\mathbf{W}^{(0)}_{\mathtt{attn}})$:
@@
where $H(p)=-\sum_{i}p_{i}\log p_{i}$ is the entropy. By setting $p_{0}$ to a uniform distribution (an unbiased baseline), the last term becomes constant and can be dropped from the loss.
Plugging Equation 6 into the compound distance term in Equation LABEL:equation:_D, we get the final form of the attention reactivation loss:
@@
This directly tackles the core goal of stabilizing convergence: it brings back suppressed attention on target tokens while keeping the updates within a stable range. Additional implementation details are provided in Appendix E.
Attack-guided velocity loss. While the attention reactivation loss brings back token-level signals, it alone cannot guarantee that the recovered attention will consistently appear in generated images. To explicitly steer the flow matching dynamics toward the target concept, we define an attack objective based on the conditional likelihood of generating an image $x$ given the target concept $c_{\text{target}}$. Our goal is to increase this likelihood by an aggressiveness factor $\eta$:
$P_{\theta+\Delta\theta}(x)\propto P_{\theta}(x)\;P_{\theta}(c_{\text{target}}\mid x)^{\eta},$ (8)
where $\theta$ are the original parameters, $\Delta\theta$ denotes the LoRA weights used for fine-tuning, and $P_{\theta}(x)$ denotes the density induced by integrating the velocity field [43] along the generative flow, playing the same role as likelihood in diffusion models. Using the identity $P_{\theta}(c_{\text{target}}\mid x)=\frac{P_{\theta}(x\mid c_{\text{target}})P_{\theta}(c_{\text{target}})}{P_{\theta}(x)}$, the gradient of the log probability $\log P_{\theta+\Delta\theta}(x)$ becomes proportional to:
$\nabla\log P_{\theta}(x)+\eta\big[\nabla\log P_{\theta}(x\mid c_{\text{target}})-\nabla\log P_{\theta}(x)\big].$ (9)
This gradient connects directly to the flow matching objective, where the model predicts a velocity field $v_{\theta}(x_{t},t)$ that governs the sample’s transport along the generative path. This yields the attack-guided velocity loss:
@@
$\begin{split}&\mathcal{L}_{\text{attack}}=\mathbb{E}\Big\|v_{\theta+\Delta\theta}(x_{t},c_{\text{target}},t)-\\ &\big[v_{\theta}(x_{t},t)+\eta\big(v_{\theta}(x_{t},c_{\text{target}},t)-v_{\theta}(x_{t},\varnothing,t)\big)\big]\Big\|_{2}^{2},\end{split}$ (10)
where $x_{t}$ denotes the latent at timestep $t$ in the denoising process, and $\varnothing$ denotes the unconditional null prompt. By optimizing this loss, the attack reshapes the generative trajectory, pulling it away from suppression and guiding the model to restore erased concepts with higher precision.
@@
LoRA consistency loss. While the previous objectives reactivate suppressed attention and amplify the target velocity, they can sometimes disturb unrelated visual elements. For instance, given the prompt “a red sports car parked beside a forest lake”, where the erased concept is “red” and unrelated content includes “sports car” and “forest lake”, our goal is to restore “red” without affecting the car or the environment layout. To achieve this, we fix a prompt $c$ that contains both the target and unrelated content, and sample 4–8 reference images using random seeds. We then train the LoRA weights $\Delta\theta$ so that the model’s predicted velocity field stays close to the original generation dynamics:
$\mathcal{L}_{\text{lora}}=\mathbb{E}_{(x_{0},x_{T},t,c)\sim\mathcal{I}_{f}}\Big\|\,v_{\theta+\Delta\theta}(x_{t},c,t)-v_{\theta}\,\Big\|_{2}^{2},$ (11)
@@
where $x_{T}$ is the Gaussian noise at the final denoising step, $x_{0}$ is the VAE-encoded latent [19] of a reference image sampled from $\mathcal{I}_{f}$, and $x_{t}=(1-t)\,x_{0}+t\,x_{T}$ is the noised latent at timestep $t$. This formulation follows Flux’s flow matching schedule, ensuring that fine-tuning stays aligned with the model’s native synthesis dynamics without affecting unrelated regions or global layout consistency.
@@
Together, these three objectives form a unified and principled framework for concept attack. The attention reactivation loss revives suppressed attention signals, the attack-guided velocity loss strengthens flow matching trajectory to reinforce the concept, and the LoRA consistency loss preserves style fidelity and protects unrelated content. By balancing precision, stability, and stealth, our method achieves robust and targeted reactivation of erased concepts.
@@ 5.1 Implementation Details
Setup. We report results on the Flux.1 [dev] [^1] https://huggingface.co/black-forest-labs/FLUX.1-dev model, an openly released distilled rectified flow transformer known for its strong prompt fidelity and high generation quality. Additional evaluations on Flux.1 [schnell] [^2] https://huggingface.co/black-forest-labs/FLUX.1-schnell and SD 3333https://huggingface.co/stabilityai/stable-diffusion-3-medium are included in Appendix H. We use the flow matching Euler sampler with 28 denoising steps. Our attack method is optimized for 200 steps using the AdamW optimizer [28] with a learning rate of 0.001 and the negative guidance factor $\eta=3$. We update 3.57MB text-related parameters ($\mathtt{add\_q\_proj}$ and $\mathtt{add\_k\_proj}$) within the dual-stream blocks. All experiments are conducted on a single NVIDIA H20 GPU (96 GB VRAM) with batch size 1.
Baselines. We evaluate our method against a range of concept erasure and attack strategies. For concept attack, we include the white-box methods UnlearnDiffAtk [57] and P4D [7], the black-box methods Ring-A-Bell and Ring-A-Bell Union [49], and the LLM reasoning-driven Reason2Attack [54]. For concept erasure, we test against leading methods including ESD [12], AC [20], EAP [3], EA [14], CP [6] and MACE [29]. More implementation details are provided in Appendix F.
Figure 3: Visual comparison of nudity and violence across attack methods under different erasure settings. Yellow framed images show original Flux.1 [dev] generations. Blue bars and blurring are added for publication purposes. Our method restores erased concepts precisely while preserving layout and unrelated content.
@@ 5.2 Results
Evaluation on NSFW concepts. NSFW concepts are widely used as benchmarks for safety evaluation. We assess our method on the I2P dataset [44], focusing on nudity and violence. For nudity, we select 109 prompts with over 50% nudity, and use NudeNet [2] for detection at a threshold of 0.6. For violence, following [49], we choose 235 prompts with over 50% inappropriateness but under 50% nudity, labeled as harmful, and apply the Q16-classifier [45] to detect violent content. Results are reported in terms of attack success rate (ASR), following [57]. Table 2 summarizes our results against major erasure methods. Our approach achieves the highest ASR across all settings, consistently outperforming existing baselines. Notably, Ring-A-Bell fails to produce effective attacks on Flux, as analyzed in Section 3. Among erasure methods, CP [6] remains the most resilient on Flux. In contrast, methods originally designed to modify cross-attention layers (e.g., AC and ESD) lose effectiveness when transferred from the U-Net structure of SD to the transformer-based architecture of Flux. This limitation is known as concept residue—the incomplete removal of erased content that makes the model vulnerable to reactivation or even amplification. As illustrated in Figure 3, our method restores erased concepts with high fidelity, recovering exposed body in nudity and blood patterns in violence, while preserving consistency in unrelated aspects (e.g., hairstyle, posture, and background layout). Other attacks, by contrast, often fail to sustain success or cause noticeable distortions in the overall composition.
Table 2: Attack evaluation on NSFW concept benchmarks. We compare concept erasure methods against multiple attack baselines. Results are reported in terms of Attack Success Rate (ASR, %; higher is better), measured on the I2P benchmark. “No Attack” indicates absence of any attack method. The performance of the original Flux.1 [dev] model is shown as a reference.
Concept Methods Flux.1 AC ESD-1 ESD-3 EA EAP CP MACE
No Attack 44.04 32.11 21.10 14.67 29.35 44.04 10.09 30.27
Ring-A-Bell 40.36 27.50 11.01 18.34 29.35 35.77 11.92 32.11
Ring-A-Bell Union 65.14 44.03 24.77 29.36 42.20 45.87 20.18 40.36
Reason2Attack 67.88 73.39 59.63 56.88 71.55 69.72 40.36 60.55
UnlearnDiffAtk 100.00 85.32 76.14 70.64 82.57 98.16 64.22 79.81
@@
No Attack 65.53 64.68 54.89 58.72 53.19 61.28 48.51 53.61
Ring-A-Bell 75.74 74.89 66.80 68.93 51.48 65.53 50.21 54.89
Ring-A-Bell Union 81.70 80.00 77.02 82.55 61.28 79.57 52.76 62.55
Reason2Attack 71.91 67.23 61.28 61.70 57.44 70.63 41.70 60.00
UnlearnDiffAtk 87.23 85.10 80.85 84.25 78.72 86.38 55.74 68.08
@@
Figure 4: Visual results on Entity, Abstraction, and Relationship. The top row shows the defense effect of MACE. Yellow framed images are the original generations from Flux.1 [dev]. UnlearnDiffAtk often produces collapsed or corrupted images (highlighted in red boxes), which are difficult for classifiers to detect, whereas our attack yields high-quality, stable generations.
Evaluation on artistic styles. We evaluate our method on two well-known painting styles: Van Gogh and Pablo Picasso, using the dataset from ConceptPrune[6] that provides 50 prompts per style. Following [57], we finetune an ImageNet-pretrained ViT-base model on WikiArt [42] and obtain a 129-class style classifier for evaluation. In reporting ASR, to avoid overly restrictive, both Top-1 and Top-3 ASR are presented, depending on whether the result is classified as the target style as the top prediction or within the top three. Table 3 shows our attack outperforms all baselines in 8 out of 12 evaluation settings, demonstrating clear and consistent improvements across artistic style benchmarks. This systematic advantage shows the effectiveness of attention reactivation and velocity control for global and highly abstract concepts, while attacks that optimize prompt embeddings are unstable and less effective on Flux.
@@
Table 4: Attack evaluation on specific category benchmarks: Entity (e.g., soccer, car), Abstraction (e.g., green, two) and Relationship (e.g., hug, kiss). CLIP classifications are reported for each category. All presented values are denoted in percentage (%).
@@
Evaluation on miscellaneousness. We evaluate our attack on 3 broader categories: Entity, Abstraction and Relationship. Here, we choose 10 concepts for each category (see Appendix F for full lists) and adopt CLIP classification as the measuring metrics. As shown in Table 4, our attack consistently outperforms baselines, with stronger gains on abstract and relational concepts, confirming its stable effectiveness across both concrete and abstract categories. Figure 4 presents the visual comparison of our method and UnlearnDiffAtk when attacking MACE erasure. Our approach stays robust even for broad or abstract concepts, where UnlearnDiffAtk may fail or produce images of poor quality.
Evaluation on efficiency. Beyond the strong performance discussed above, our attack is also highly efficient. While UnlearnDiffAtk and P4D rely on costly PGD optimization for each prompt (exceeding 20 minutes per prompt), our method fine-tunes once and can be directly applied to all related prompts of the same concept. After fine-tuning, attacking a single prompt requires only standard image generation time (about 0.5 minutes), demonstrating both practicality and scalability.
Figure 5: Layer-wise attack gains across dual stream blocks. Attack gain is defined as the relative increase in activation strength of our attack method compared to the erased model (EA).
Evaluation on layout consistency. Table 5 reports average consistency between images before and after attacking EA, evaluated by LPIPS [56] and CLIP similarity. Our method achieves the best balance between strong attack and minimal layout distortion, maintaining both visual coherence and fidelity across all concept categories.
Table 5: Visual consistency between erased (EA) and attacked images measured by LPIPS (perceptual differences) and CLIP similarity (semantic similarity). Lower LPIPS and higher CLIP indicate better preservation of layout and unrelated visual content.
Nudity Violence Style Entity Abstract Relation
@@
Ring-A-Bell 0.21 0.82 0.55 0.73 0.82 0.88 0.47 0.89 0.69 0.83 0.43 0.74
P4D - - - - 0.73 0.76 - - - - - -
Reason2Attack 0.58 0.71 0.87 0.68 - - - - - - - -
UnlearnDiffAtk 0.83 0.55 0.73 0.69 0.71 0.65 0.66 0.72 0.82 0.81 0.77 0.82
[…]
Evaluation on layer-wise gains. To understand where our attack revives erased signals, we inspect Flux’s 19 dual-stream blocks. As shown in Figure 5, different concepts reappear at different stages: “nudity” resurfaces early and again at the final block, “Van Gogh” shows clear boosts in the middle–late layers, and “soccer” climbs more gradually but stays consistently positive. This suggest that the attack taps into specific representation stages rather than pushing all layers uniformly; more discussions are in Appendix G.
@@
Ablation study. To evaluate the effectiveness of our loss functions, we conduct an ablation study on celebrities. We attack the EA erasure using a subset of CelebA [27], containing 50 celebrities for attack and 50 for retention. Different variations are quantified in Table 6. The velocity objective $\mathcal{L}_{\text{attack}}$ sets the groundwork by steering the model away from suppression, making concept reactivation reliable across settings. The attention reactivation term $\mathcal{L}_{\text{attn}}$ sharpens targeting by boosting suppressed attention, significantly improving Acce. As shown in Figure 6, its L2 regularizer constrains activation strength for reliable concept revival, and entropy maintains style and visual coherence. Additionally, the consistency term $\mathcal{L}_{\text{lora}}$ further anchors background and irrelevant attributes, keeping details such as clothing, hairstyle, and posture consistent between the erased and attacked output, serving as a key contributor to high Accir. When integrated, these components form a synergistic objective: attention reactivation drives precision, velocity control enhances robustness, and LoRA regularization safeguards fidelity, together delivering the strongest and most balanced reactivation of erased concepts.
@@
User study and others. To evaluate human perception of our method’s effectiveness, we conduct a four-aspect user study with 20 non-artist participants, each contributing around 200 responses. As shown in Figure 7, our method performs strongly across all aspects, reflecting its well-rounded capability in concept attack. Appendix H provides more details, including additional results, extended ablations, dataset details, pretrained classifier specifications, further visualizations and VLM assessments.
Figure 7: User Study. We design an interface (details in Appendix H) that presents AIGC results under different attacks and ask users to rate them from 1–5. Across all four evaluation dimensions, ReFlux achieves the strongest overall performance.